Lin Corporation has a single product whose selling price is $120 per unit and whose variable expense is $80 per unit. The company’s monthly fixed expense is $50,000. Required: 1. Calculate the unit sales needed to attain a target profit of $10,000. 2. Calculate the dollar sales needed to attain a target profit of $15,000. I mostly need help with finding the cm ratio for #2. Showing work for all steps would be helpful! Thanks (For all requirements, do not round intermediate calculations.)
Answer #1
1
Unit sales needed = (10000+50000)/(120-80)= 1500
2
CM ratio = (120-80)/120=33.33%
Dollar sales needed =(15000+50000)/33.33%= 195000